Student123:

theodore know hat 0.271 =271/100 he believes that this is read as two hundredths seventy-one tenths is he correct

Student123:

help

supie:

Ok so you know that the first number after the decimal is the tenths and the second one is hundreths and the 3rd one is thousandths and so on right?

Student123:

yed i know that

Student123:

yes

supie:

So then do you think he is correct?

Student123:

i dont know

supie:

Ok, lemme explain, 

supie:

Now, for example, 0.738 would be written as `seven hundred and 8 thousandths` since it is in the thousandths

supie:

So what do you think now?

supie:

(We can ignore the part of the problem that says `0.271 ="271/100",` that is just saying that 0.271 can be written as a fraction, and if it was written as a fraction, that is what it would look like). The only part we need is `0.271.`

Student123:

so i think he's incorret

supie:

Yes, he is incorrect, the correct way for him to write it would be `two hundred seventy-one thousandths`

supie:

Notice he said `tenths` and mine says `thousandths`.

Student123:

so he's correct

supie:

No, I just said notice he said `tenths` and I said `thousandths`

Student123:

ok

supie:

So you were right! Good luck!

Student123:

now i need help with something else

supie:

Make a new post please, and close this one.

Student123:

ok

jhonyy9:

@supie pardon for my interfer but i dont understand it 271/100 = 2,71 this is what you said above ? i dont see it in this way

supie:

No, 0.271.

supie:

We are trying to find what 0.271 can be written as.

jhonyy9:

theodore know hat 0.271 =271/100 he believes that this is read as two hundredths seventy-one tenths is he correct 0,271 not equal 271/100

jhonyy9:

0,271 = 271/1000

supie:

Oh yeah....

supie:

I didn't write the problem though...he did...

jhonyy9:

and now how will be ??? @supie

supie:

????

jhonyy9:

please correct the way what you ve wrote above

jhonyy9:

@Laylalyssa your opinion pls ?

supie:

It's a mistake in the question. He wrote it himself... he messed up on it I think @student123

jhonyy9:

@Student123 pls come back this is not solved correct

jhonyy9:

theodore know hat 0.271 =271/100 he believes that this is read as two hundredths seventy-one tenths is he correct

jhonyy9:

this is the posted problem

jhonyy9:

@Laylalyssa your opinion about this pls. ?

jhonyy9:

my opinion that @Student123 missed one zero from denominator bc. 0,271 = 271 /1000

jhonyy9:

this in the first step

supie:

That's what I thought too, so does that mean I was right before?

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @Laylalyssa your opinion about this pls. ? \(\color{#0cbb34}{\text{End of Quote}}\) it does mean he was right b4 bc 0.271 is two hundredths and 71 tenths also what supie sed basically the same thing

jhonyy9:

i think no - bc. 0,271 is 2 tenth 7 hundredths 1 thousandths

supie:

It's a decimal `.` not a comma `,` yk?

jhonyy9:

0.271 is two hundred seventy-one thousandths

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 0.271 is two hundred seventy-one thousandths \(\color{#0cbb34}{\text{End of Quote}}\) That's what I said.

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Yes, he is incorrect, the correct way for him to write it would be `two hundred seventy-one thousandths` \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

2 mean there thenths

jhonyy9:

0,2 is two tenths

Laylalyssa:

so its 0.271 2 tenths 7 hundredths 1 thousandths ?

jhonyy9:

yes and together how you exprime it ?

Laylalyssa:

exprime? is that even a word, I have never heard of it..

Laylalyssa:

or is it express

jhonyy9:

express yes - sorry

jhonyy9:

exprime is french

Laylalyssa:

oh ya I dont speak French anyways what do you mean by express the question was asking if the person was correct or not right?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa so its 0.271 2 tenths 7 hundredths 1 thousandths ? \(\color{#0cbb34}{\text{End of Quote}}\) this i ve thought ask you how you express it in a word ?

supie:

We are trying to write 0.271 in words Since there is three numbers after the decimal, then we would say thousandths. Since the number is `271` then we would say two hundred-and-one thousandths. Yes?

Laylalyssa:

I think yes

jhonyy9:

and the 7 where is ?

Laylalyssa:

u mean where is the seven?

jhonyy9:

say two hundred-and-one thousandths.

jhonyy9:

from there

supie:

I said `two hundred-and-one thousandths ` but I meant `two hundred seventy-one thousandths`

jhonyy9:

yes in this way is now correct wrote

Laylalyssa:

so thats it

supie:

That's what I said to begin with though.

Laylalyssa:

I like how we solved the problem and the kid not even here smh

jhonyy9:

yes you see it right - the kid was go away

jhonyy9:

dont make interest in ...

Laylalyssa:

ya,